Problem 3: A 7200-kg rocket blasts off vertically from the launch pad with a constant upward acceleration of 2.20m/s? and feels no appreciable air resistance. Whe it has reached a height of 555 m, its engines suddenly fail; the only force acting on it is now gravity. a. What is the maximum height this rocket will reach above the launch pad? b. How much time will elapse after engine failure before the rocket comes crashing down to the launch pad, and how fast will it be moving just before it crashes? c. Sketch av-t, uy-t, and y-t graphs of the rockets motion from the instant of blast- off to the instant just before it strikes the launch pad.
Answer #1

a. during upward acceleration,

d1 = 555 m

and d = v0 t + a t^2 / 2

555 = 0 + (2.2) t^2 /2

t1 = 22.462 sec

v1 = v0 + a t = 0 + (2.2)(22.462) = 49.42 m/s


during free fall,

vf^2 - vi^2 = 2 a d

0^2 - 49.42^2 = 2(-9.8) d2

d2 = 124.6 m

(A) H_max = d1 +d2 = 679.6 m


(B) t1 = 22.462 sec  

0 = 49.42 - 9.8 t2

t2 = 5.04 s

  
after that,

- 679.6 = 0 - 9.8 t3^2 / 2

t3 = 11.78 sec  


t = t1 + t2 + t3 = 39.3 sec .......Ans

v = 0 + (9.8 x 11.78) = 115.4 m/s ,,,, Ans